OpenStudy (anonymous):

How do i solve this initial value problem:

OpenStudy (anonymous):

and where's the problem?

OpenStudy (anonymous):

\[ds/dt=8\sin^2(t+(\pi/2)), s(0)=8\]

OpenStudy (anonymous):

can someone just tell, what it is we are trying to do here in this problem

OpenStudy (anonymous):

are we integrating to find F or what? i dont get

OpenStudy (zarkon):

\[ds/dt=8\sin^2(t+(\pi/2))\] \[ds=8\sin^2(t+(\pi/2))dt\] \[\int ds=\int8\sin^2(t+(\pi/2))dt\] \[=\cdots\]

OpenStudy (anonymous):

now, when integratins ds, does that just intergrate to s

OpenStudy (zarkon):

yes \[\int ds=s+c\]

OpenStudy (anonymous):

can i ask you one more thing...

OpenStudy (anonymous):

I had this problem : \[6+\int\limits_{a}^{x}f(t)/t^2 dt=2\sqrt{x}\]

OpenStudy (anonymous):

i was told : for all x>0 find the function f and a number a such that

OpenStudy (zarkon):

if you need to find f...then use the fundamental theorem of calculus and differentiate both sides..then solve for f

OpenStudy (anonymous):

are we trying to isolate f(t)

OpenStudy (zarkon):

f(x)

OpenStudy (anonymous):

now what about this number "a", what is taht suppose to be?

OpenStudy (zarkon):

\[\frac{d}{dx}\int\limits_{a}^{g(x)}f(t)dt=f(g(x))g'(x)\]

OpenStudy (anonymous):

oh i see a is suppose to be like the lower limit of integration of something like that

OpenStudy (zarkon):

yes...so it disappears once you differentiate.

OpenStudy (anonymous):

how would i go about solving for that?

OpenStudy (zarkon):

you need to differentiate both sides of your equation

OpenStudy (anonymous):

i got : f(x)=x^2/x^1/2

OpenStudy (anonymous):

oh then just plug a into that

OpenStudy (zarkon):

the a plays no role in this problem

OpenStudy (anonymous):

okay, but i was told to find a number "a"

OpenStudy (zarkon):

find a number a such that what ?

OpenStudy (anonymous):

such that: \[6+\int\limits_{a}^{x}f(t)/t^2dt=2\sqrt{x}\]

OpenStudy (zarkon):

ok...jas

OpenStudy (anonymous):

what?

OpenStudy (zarkon):

jas...just a sec

OpenStudy (anonymous):

okay

OpenStudy (zarkon):

look like 9

OpenStudy (zarkon):

replace \[f(x)=x^{3/2}\] into the equation then integrate

OpenStudy (anonymous):

what r u saying? like 6+x^3/2=2sqrtx

OpenStudy (zarkon):

\[6+\int\limits_{a}^{x}\frac{t^{3/2}}{t^2} dt=2\sqrt{x}\]

OpenStudy (zarkon):

\[6+\int\limits_{a}^{x}t^{-1/2} dt=2\sqrt{x}\]

OpenStudy (zarkon):

\[6+\left.2t^{1/2}\right|_a^{x}=2\sqrt{x}\] \[6+2x^{1/2}-2a^{1/2}=2\sqrt{x}\] \[6+2\sqrt{x}-2\sqrt{a}=2\sqrt{x}\]

OpenStudy (zarkon):

a=9

OpenStudy (anonymous):

no more questions from me after this one, why is it that we find the "a" this way?

OpenStudy (zarkon):

how else would we find it?

OpenStudy (anonymous):

but, i mean, why is this the way that we find it?

OpenStudy (zarkon):

we need to isolate a. if a is stuck on the integral sign I can't isolate it.

OpenStudy (anonymous):

thanks Zarkon, you the best, say hello to buzz for me

OpenStudy (anonymous):

you know from toystory

OpenStudy (zarkon):

Zarkon is from Voltron ;)